# Thermodynamic Cost of Regeneration in a Quantum Stirling Cycle

### v1

@@ 3 Conclusion
We investigated the regenerative quantum Stirling cycle within the standard weak-coupling, Markovian open quantum system framework. Inspired by the classical regenerator model, to date, virtually all treatments model the regenerator as a passive heat buffer that stores heat during the isochoric stroke (2) and returns it in stroke (4), with the corresponding work cost typically left implicit. In the reduced open-system description, the regenerator is treated as an effective heat reservoir that brings the working substance to thermal equilibrium at the end of the isochoric strokes. When this idealized, cost-free regenerator picture is adopted, the resulting bookkeeping may suggest very high apparent efficiencies, in some cases even exceeding the Carnot bound. In this work, based on the weak-coupling open quantum system description of the system-bath interactions, we point out that the operation of the regenerator implicitly requires the heat stored at a low temperature level to be made usable at a higher temperature level, i.e., that the stored (recovered) heat be pumped from the lower temperature to the higher temperature. By the second law of thermodynamics, this process necessarily requires an additional work input. We then accounted for this hidden work input as an explicit regeneration cost within the working-substance thermodynamic description, and we used it to redefine the cycle efficiency by taking the net work produced by the working substance per unit total energy input. We revisited the regenerative Stirling cycle using two working substances: a single spin-$1/2$ and a pair of interacting spin-$1/2$ particles. In both cases, we re-evaluated the cycle performance using our modified efficiency, which includes the regeneration cost taken at its minimum value set by the Carnot heat pump limit. For comparison, we also analyzed the conventional Stirling cycle without regeneration and computed its efficiency under the same conditions, in order to assess whether the regenerator still provides an advantage despite this cost. Our results demonstrate that the modified efficiency including $W_{\mathrm{cost}}$ remains below the Carnot efficiency, while still being higher than the efficiency of the conventional Stirling cycle. For the conventional Stirling cycle, we proved that the Carnot efficiency is a rigorous upper bound by using the quantum relative entropy formulation. In contrast, for the regenerative Stirling cycle within our phenomenological reduced description, we showed that the Carnot efficiency can be guaranteed as an upper bound provided that the regeneration cost $W_{\mathrm{cost}}$ exceeds a sufficient lower bound. Finally, to motivate future work, we proposed three possible quantum models in which the regenerator is treated as an active component of the cycle, and we briefly discussed the physical motivation behind each proposal.
@@
Eq. (14) is the simplest expression that can be given for the Carnot deficit of the regenerative cycle. The first term on the right-hand side is always non-negative. On the other hand, when we take $W_{\mathrm{cost}}$ to be its lower bound determined by the Carnot heat pump limit, the sign of the second term—and whether it is smaller or larger than the first term—remains uncertain. Therefore, we cannot prove that $\eta_{C}-\eta\geq 0$. Instead, using Eq. (14), we can derive a sufficient lower bound on $W_{\mathrm{cost}}$ that guarantees $\eta_{C}-\eta\geq 0$. Since $(Q_{h}+W_{\mathrm{cost}})>0$, we obtain a sufficient lower bound on $W_{\mathrm{cost}}$ ensuring that the numerator of Eq. (14) is non-negative:

### v2

@@ 3 Conclusion
We investigated the standard four-stroke regenerative quantum Stirling cycle, which assumes local thermal equilibrium at each stage, within the standard weak-coupling, Markovian open quantum system framework. Inspired by the classical regenerator model, to date, virtually all treatments model the regenerator as a passive heat buffer that stores heat during the isochoric stroke (2) and returns it in stroke (4), with the corresponding work cost typically left implicit. In the reduced open-system description, the regenerator is treated as an effective heat reservoir that brings the working substance to thermal equilibrium at the end of the isochoric strokes. When this idealized, cost-free regenerator picture is adopted, the resulting bookkeeping may suggest very high apparent efficiencies, in some cases even exceeding the Carnot bound. In this work, based on the weak-coupling open quantum system description of the system-bath interactions, we point out that the operation of the regenerator implicitly requires the heat stored at a low temperature level to be made usable at a higher temperature level, i.e., that the stored (recovered) heat be pumped from the lower temperature to the higher temperature. By the second law of thermodynamics, this process necessarily requires an additional work input. We then accounted for this hidden work input as an explicit regeneration cost within the working-substance thermodynamic description, and we used it to redefine the cycle efficiency by taking the net work produced by the working substance per unit total energy input. We revisited the regenerative Stirling cycle using two working substances: a single spin-$1/2$ and a pair of interacting spin-$1/2$ particles. In both cases, we re-evaluated the cycle performance using our modified efficiency, which includes the regeneration cost taken at its minimum value set by the Carnot heat pump limit. For comparison, we also analyzed the conventional Stirling cycle without regeneration and computed its efficiency under the same conditions, in order to assess whether the regenerator still provides an advantage despite this cost. Our results demonstrate that the modified efficiency including $W_{\mathrm{cost}}$ remains below the Carnot efficiency, while still being higher than the efficiency of the conventional Stirling cycle. For the conventional Stirling cycle, we proved that the Carnot efficiency is a rigorous upper bound by using the quantum relative entropy formulation. In contrast, for the regenerative Stirling cycle within our phenomenological reduced description, we showed that the Carnot efficiency can be guaranteed as an upper bound provided that the regeneration cost $W_{\mathrm{cost}}$ exceeds a sufficient lower bound. Finally, to motivate future work, we proposed three possible quantum models in which the regenerator is treated as an active component of the cycle, and we briefly discussed the physical motivation behind each proposal.
@@
Eq. (14) is the simplest expression that can be given for the Carnot deficit of the regenerative cycle. The first term on the right-hand side is always non-negative. On the other hand, under general conditions, the sign of the second term and its exact relationship with the first term remain analytically uncertain. When we evaluate this expression by setting $W_{\mathrm{cost}}$ to its theoretical lower bound determined by the Carnot heat pump limit, i.e., $W_{\mathrm{cost}}=|Q_{2}|\,((T_{h}-T_{c})/T_{c})$, the sign of the second term is unconditionally non-negative for temperature regimes where $T_{h}\geq 2T_{c}$. In this regime, we therefore always obtain $\eta_{C}>\eta$. However, in the regime where $T_{h}<2T_{c}$, this second term can become negative. Crucially, a negative second term does not imply that the Carnot deficit ($\eta_{C}-\eta$) becomes negative. The first term in the equation remains positive and generally dominates the second term, preserving the condition $\eta_{C}-\eta>0$. This behavior is explicitly demonstrated by the exact numerical results presented in Figs. 2 and 3. In these plots, the bath temperatures are set to $T_{h}=3$ and $T_{c}=2$, which precisely falls into the $T_{h}<2T_{c}$ regime. As seen in the figures, even in this analytically ambiguous regime, the evaluated efficiency consistently stays strictly below the Carnot limit (see red dashed lines). Consequently, the absence of a closed-form analytical proof in this regime does not signal a violation of the second law, but rather reflects the mathematical limitation of applying a reversible macroscopic bound to an irreversible reduced quantum process. Instead of attempting to force a general proof from the Carnot limit, we use Eq. (14) to provide a sufficient lower bound on $W_{\mathrm{cost}}$ that mathematically guarantees $\eta_{C}-\eta\geq 0$ across all regimes. Since $(Q_{h}+W_{\mathrm{cost}})>0$, we obtain a sufficient lower bound on $W_{\mathrm{cost}}$ ensuring that the numerator of Eq. (14) is non-negative: